What mass of HClO4 should be present in 0.600 L of solution to obtain a solution with each of the following pH values?

Part A pH= 2.70 Express your answer using two significant figures. m = _______ g

Part B pH= 1.60 Express your answer using two significant figures. m = _______ g  

Part C pH= 0.60 Express your answer using two significant figures. m = ______ g
Answer #1

A)

use:

pH = -log [H+]

2.7 = -log [H+]

[H+] = 2.0*10^-3 M

So, concentration of HClO4 is 2.0*10^-3 M

volume , V = 0.6 L

use:

number of mol,

n = Molarity * Volume

= 2*10^-3*0.6

= 1.2*10^-3 mol

Molar mass of HClO4,

MM = 1*MM(H) + 1*MM(Cl) + 4*MM(O)

= 1*1.008 + 1*35.45 + 4*16.0

= 100.458 g/mol

use:

mass of HClO4,

m = number of mol * molar mass

= 1.2*10^-3 mol * 1.005*10^2 g/mol

= 0.1205 g

Answer: 0.12 g

B)

use:

pH = -log [H+]

1.6 = -log [H+]

[H+] = 2.5*10^-2 M

So, concentration of HClO4 is 2.5*10^-2 M

volume , V = 0.6 L

use:

number of mol,

n = Molarity * Volume

= 2.5*10^-2*0.6

= 1.5*10^-2 mol

Molar mass of HClO4,

MM = 1*MM(H) + 1*MM(Cl) + 4*MM(O)

= 1*1.008 + 1*35.45 + 4*16.0

= 100.458 g/mol

use:

mass of HClO4,

m = number of mol * molar mass

= 1.5*10^-2 mol * 1.005*10^2 g/mol

= 1.507 g

Answer: 1.5 g

C)

use:

pH = -log [H+]

0.6 = -log [H+]

[H+] = 0.2512 M

So, concentration of HClO4 is 0.2512 M

volume , V = 0.6 L

use:

number of mol,

n = Molarity * Volume

= 0.2512*0.6

= 0.1507 mol

Molar mass of HClO4,

MM = 1*MM(H) + 1*MM(Cl) + 4*MM(O)

= 1*1.008 + 1*35.45 + 4*16.0

= 100.458 g/mol

use:

mass of HClO4,

m = number of mol * molar mass

= 0.1507 mol * 1.005*10^2 g/mol

= 15.14 g

Answer: 15 g